# Multi-task neural diffusion processes for uncertainty-quantified wind power prediction

### v1

@@ 4.2 Noise model architecture
The noise model architecture follows [22], designed to generate sensible prior distributions over functions. While any NN could in principle serve as the denoiser, certain structural properties are desirable: There are two key requirements:
  • (i)
    Input-size agnosticity: the model should work for arbitrary numbers of input points $N$ and input dimensions $D$, producing consistent priors whether inference is carried out at one or many locations. This is achieved by replicating the outputs $y_{t}\in\mathbb{R}^{N}$ across $D$ copies so that the network weights do not depend directly on $N$ or $D$. As a result, one trained NDP can be applied across datasets of varying size and dimensionality.
  • (ii)
    Equivariance and invariance: predictions should not depend on the ordering of inputs, context points, or feature dimensions. Reordering inputs should not change the probability of the data, just as kernels such as the RBF in GPs are invariant to permutations. To enforce the equivariance and invariance, [22] proposed a bi-dimensional attention block, $A_{t}:\mathbb{R}^{N\times D\times H}\to\mathbb{R}^{N\times D\times H}$. This multi-head self-attention mechanism operates jointly across input locations and feature dimensions. Attention weights are computed via dot-product similarity, and multiple heads are run in parallel to capture complementary relationships. This design ensures permutation equivariance and invariance while retaining the flexibility of neural architectures, addressing limitations that are often overlooked in standard NN models.
4.3 Multi-task neural diffusion processes
We propose a novel architectural addition to extend NDPs to MT-NDPs, borrowing key items from [18]. The broad idea is to add a task encoder that can take context points and use that information to identify the task the model is performing a prediction on. This information is represented in the encoding space by a single vector $v\in\mathbb{R}^{\kappa}$. That is, the encoding space has dimension $\kappa$, allowing enough flexibility for the model to distinguish between different tasks, i.e., wind turbines. The dimension $\kappa$ is assumed to be small; for example, in our application to SCADA data described in Section 5, we set $\kappa=8$. This vector can then be propagated downstream to the diffusion architecture, which remains unchanged. The way we pass the information downstream is simply to concatenate these extra $\kappa$ dimensions onto the other input features. These $\kappa$ additional features will be the same for all points in the same training (or testing) sample.
4.3.1 Forward Process
@@ 4.3.2 Reverse Process
Therefore, given the initial input locations and targets $T=(x_{0}\in\mathbb{R}^{N\times D},y_{0}\in\mathbb{R}^{N})$ and context points $C=(x^{c}_{0}\in\mathbb{R}^{M\times D},y^{c}_{0}\in\mathbb{R}^{M})$, after the encoder we preserve $T$ and are left with $N$ copies of a vector $\bar{v}\in\mathbb{R}^{\kappa}$ that we concatenate to the other input features of $x_{o}$ to leave us with $D+\kappa$ dimensional input features, and a modified set of input locations and targets: $T^{*}=(x^{*}_{0}\in\mathbb{R}^{N\times(D+\kappa)},y^{*}_{0}\in\mathbb{R}^{N})$ that is then passed through the NDP de-noising model $\epsilon^{NDP}_{\theta}(x^{*}_{t},y^{*}_{t},t):\mathbb{R}^{N\times(D+\kappa)}\times\mathbb{R}^{N}\to\;\mathbb{R}^{N}$.
By constructing the task encoder in this way, we ensure agnosticity to input size. During training, the number of context points in each set is sampled uniformly from the integers 0 to $50$. This setting exposes the model to situations with no context, with few context points, and with many context points. As a result, the model learns both a global prior when no context is provided and task-specific priors when context points are available.
Equivariance is preserved by taking the mean $\bar{v}\in\mathbb{R}^{\kappa}$, which ensures that the mapping is invariant to the ordering of context points. The target points remain unchanged by the encoder. Invariance, however, depends on the encoder architecture. Here we use a simple MLP, which does not guarantee invariance. Achieving this property would require multi-head self-attention blocks, as in the NDP architecture, but this would substantially increase model complexity and computational cost. In the specific application to wind farms, where the ordering of input features can be fixed across training and testing, the lack of invariance is less problematic than it would be in broader meta-learning settings. Details of hidden layer sizes are provided in Section 5.2.
4.3.3 Objective
@@
4.3.4 Sampling
Unconditional sampling remains unchanged from the NDP setup described in Section 4.1.4, as we use no context points at all. Simply take $y_{T}\sim\mathcal{N}(0,I)$, and then for $t=T,\dots,1$ apply the parametrised kernel. For conditional sampling, we still follow the process for conditional sampling of NDPs described in Section 4.1.4, but now using our upgraded de-noising model $\epsilon^{MT}_{\theta}$:
@@ 4.3.4 Sampling
5 Experimental design
We evaluate the proposed models on wind power prediction tasks using SCADA data from the Kelmarsh wind farm. As the number of input variables and model complexity increase, training times grow substantially. Inference with NDPs can also be slow, since both conditional and unconditional sampling require $T=500$ diffusion steps, each involving a neural network evaluation. By contrast, GPs offer fast inference for small datasets, but quickly become intractable for larger datasets. Sparse GPs have been studied on the same dataset [15], but suffer from underconfidence in predictive uncertainty.
A key advantage of NDPs and related meta-learning models is their ability to adapt to unseen tasks with only a small number of context points. To test this property, we adopt a challenging setup where models are trained on a subset of turbines and evaluated on a previously unseen turbine. Context points, drawn from the training set of the test turbine, guide the model during inference. GPs do not use context sets explicitly; all training data acts as context.
@@ 5.1 Training and testing framework
All features are standardised per wind turbine, using the mean and standard deviation from the wind turbine’s training set. This step is necessary for both NDPs and GPs. To ensure fair comparisons, we use the same test functions across all models and context sizes. Context sets of size 0, $25$, and $50$ points are considered. Each evaluation uses $30$ test functions of length $100$.
5.2 Model paramemeters
For all NDP variants, we use a cosine noise schedule [24] with $T=500$ diffusion steps. The denoising neural network has $4$ hidden layers of dimension $64$, with multi-head self-attention using $\kappa=8$ heads. The optimiser is ADAM [26], following the configuration of [22], with a $20$-epoch warmup, $200$-epoch decay, learning rate schedule ($2\times 10^{-5}\to 10^{-3}\to 10^{-5}$), and an EMA weight of $0.995$. For sampling, we generate $200$ reverse diffusion trajectories in the one-dimensional case and $100$ in higher-dimensional settings, balancing predictive sharpness with computational cost.
In the multi-task framework, the task encoder is an MLP with four $4$ layers of size $64$, GeLU activations, and an $8$-dimensional output embedding. The mean embedding across all context points forms a task representation vector. Training is performed for $250$ epochs using batches of $32$ function samples of length $100$, with $500$ samples drawn per wind turbine ($5$ tasks/turbines in total), and $100$ randomly shuffled function samples presented per epoch.
@@ 5.3 Model evaluation
To assess point prediction accuracy, we compute the following per-sample losses for each test set: (i) mean absolute error (MAE), defined as $\mathrm{MAE}s=\tfrac{1}{N}\sum{i=1}^{N}\lvert y_{i}-\hat{y}_{i}\rvert$ and (ii) root mean squared error (RMSE), defined as $\mathrm{RMSE}s=\sqrt{\tfrac{1}{N}\sum{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$.
@@ 6 Results
We first compare a one-dimensional GP with a one-dimensional NDP, using only wind speed as a feature. In this setting, wind turbines are nearly indistinguishable, consistent with exploratory analysis (Fig. 1(a)). For the GP, we mimic the NDP setup, training on turbines 2266 and testing on turbine $1$. Due to scalability limits, the GP is trained on a subsample of $2000$ points evenly drawn from the five turbines. We compare this GP baseline only to the NDP with no context points, ensuring a fair assessment of prior quality.
We next compare a one-dimensional NDP, using wind speed as a feature, with a five-dimensional NDP. The additional inputs are wind direction, encoded as its sine and cosine components, nacelle temperature, and transformer temperature. These variables were selected based on exploratory analysis shown in Fig. 2, which revealed strong or nonlinear relationships with wind power output.
@@
6.1 One-dimensional GP against one-dimensional NDP
We first compare the one-dimensional NDP without context points against a GP baseline to ensure a fair evaluation. The effect of context on NDP performance is examined separately in Sections 6.2 and 6.4.
Figure 6: Fit of Gaussian process (GP) (left) and neural diffusion process (NDP) (right) on a test sample of 100 points from turbine 1, with 95% confidence intervals (CI) on normalised wind speed and power output. 1D = One-dimensional.
@@ 6.1 One-dimensional GP against one-dimensional NDP
(a) Mean absolute error (MAE, left) and root mean square error (RMSE, right) distributions for Gaussian process (GP) and neural diffusion process (NDP) over $30$ test samples, each with $100$ points (turbine 11).
(b) Observed vs. theoretical coverage probabilities for Gaussian process (GP) and neural diffusion process (NDP) over $30$ test samples, each with $100$ points (turbine 11).
Figure 7: Comparison of Gaussian process (GP) and neural diffusion process (NDP) on turbine 11 test data. (a) Point prediction accuracy (MAE and RMSE distributions). (b) Predictive uncertainty calibration (coverage probabilities). 1D = One-dimensional, ST= single-task
Considering all $30$ test samples of $100$ points each, both models achieve comparable point prediction accuracy. As shown in Fig. 7(a), the distributions of MAE and RMSE for the GP and NDP are nearly indistinguishable, indicating that both capture the mean power–wind speed relationship effectively.
The distinction between the models becomes clear when evaluating predictive uncertainty. Fig. 7(b) shows the empirical coverage against the nominal coverage level. A perfectly calibrated model would follow the diagonal $y=x$ line. While the NDP is not flawless, it consistently lies closer to this ideal than the GP, which suffers from pronounced underconfidence across coverage levels. This highlights a key strength of NDPs: the ability to generate well-calibrated uncertainty estimates without relying on Gaussian assumptions.
6.2 One-dimensional NDP against five-dimensional NDP
Figure 8: Distributions of mean absolute error (MAE, left) and root mean square error (RMSE, right) for the one-dimensional (1D) and five-dimensional (5D) neural diffusion process (NDP) models, evaluated with 0, $25$, and $50$ context points. Results are aggregated over $30$ test functions of $100$ points each from turbine $1$. ST = single-task.
Figure 9: Observed vs. theoretical coverage probabilities for the one-dimensional (1D) and five-dimensional (5D) neural diffusion process (NDP) models with 0, $25$, and $50$ context points. Results are averaged over $30$ test functions of $100$ points each from turbine $1$. ST = single-task.
6.3 Aggregated results
Table 1: Out-of-sample performance of Gaussian process (GP) and single-task neural diffusion process (ST-NDP) models on turbine $1$, averaged over $30$ test functions ($100$ points each). Metrics reported: mean absolute error (MAE, kW), root mean squared error (RMSE, kW), and coverage error (CE, %).
Model # of features Context MAE (kW) RMSE (kW) CE (%)
GP 11 0 $51.636$ $75.596$ $18.075$
ST-NDP 11 0 $52.603$ $75.139$ $4.584$
$25$ $51.273$ $74.568$ $4.181$
$50$ $50.545$ $74.137$ $3.938$
ST-NDP 33 0 $60.967$ $88.096$ $9.731$
$25$ $52.338$ $74.312$ $6.750$
$50$ $50.855$ $73.537$ $5.369$
ST-NDP 55 0 $49.429$ $67.685$ $8.351$
$25$ $46.962$ $66.007$ $6.029$
$50$ $45.791$ $65.301$ $4.816$
@@ 6.4.1 Encoder analysis
These findings have two key implications for evaluation. First, turbine $1$ is not an informative test case for the MT-NDP: because its embeddings remain close to the global mean, the model receives little benefit from task-specific adaptation, and its predictions are nearly identical to those of the ST-NDP. Second, turbine $6$ stands out as the most challenging and distinctive case in the wind farm, forcing the model to deviate from the global prior and rely heavily on context information for effective adaptation.
For this reason, in the following sections, we focus our quantitative evaluation on turbine $6$. This choice provides a stronger and more discriminative test of the MT-NDP’s ability to capture turbine-specific behaviour, highlighting the advantages of multi-task learning in settings where tasks differ substantially.

### v2

@@ 2.1 Noise model architecture
The noise model architecture follows [dutordoir2023neural], designed to generate sensible prior distributions over functions. While any NN could in principle serve as the denoiser, certain structural properties are desirable: There are two key requirements: (i) Input-size agnosticity: the model should work for arbitrary numbers of input points $N$ and input dimensions $D$, producing consistent priors whether inference is carried out at one or many locations. This is achieved by replicating the outputs $y_{t}\in\mathbb{R}^{N}$ across $D$ copies so that the network weights do not depend directly on $N$ or $D$. As a result, one trained NDP can be applied across datasets of varying size and dimensionality. (ii)Equivariance and invariance: predictions should not depend on the ordering of inputs, context points, or feature dimensions. Reordering inputs should not change the probability of the data, just as kernels such as the radial basis function (RBF) in GPs are invariant to permutations. To enforce the equivariance and invariance, [dutordoir2023neural] proposed a bi-dimensional attention block, $A_{t}:\mathbb{R}^{N\times D\times H}\to\mathbb{R}^{N\times D\times H}$. This multi-head self-attention mechanism operates jointly across input locations and feature dimensions. Attention weights are computed via dot-product similarity, and multiple heads are run in parallel to capture complementary relationships. This design ensures permutation equivariance and invariance while retaining the flexibility of neural architectures, addressing limitations that are often overlooked in standard NN models.
3 Multi-task neural diffusion processes (MT-NDPs)
We propose a novel architectural addition to extend NDPs to MT-NDPs, borrowing key items from [kim2022multitaskneuralprocesses]. The broad idea is to add a task encoder that can take context points and use that information to identify the task the model is performing a prediction on. This information is represented in the encoding space by a single vector $v\in\mathbb{R}^{\kappa}$. That is, the encoding space has dimension $\kappa$, allowing enough flexibility for the model to distinguish between different tasks. The dimension $\kappa$ is assumed to be small; for example, in our experiments described in Section 4, we set $\kappa=8$. This vector can then be propagated downstream to the diffusion architecture, which remains unchanged. The way we pass the information downstream is simply to concatenate these extra $\kappa$ dimensions onto the other input features. These $\kappa$ additional features will be the same for all points in the same training (or testing) sample.
3.1 Forward Process
@@ 3.2 Reverse Process
Therefore, given the initial input locations and targets $T=(x_{0}\in\mathbb{R}^{N\times D},y_{0}\in\mathbb{R}^{N})$ and context points $C=(x^{c}_{0}\in\mathbb{R}^{M\times D},y^{c}_{0}\in\mathbb{R}^{M})$, after the encoder we preserve $T$ and are left with $N$ copies of a vector $\bar{v}\in\mathbb{R}^{\kappa}$ that we concatenate to the other input features of $x_{0}$ to leave us with $D+\kappa$ dimensional input features, and a modified set of input locations and targets: $T^{*}=(x^{*}_{0}\in\mathbb{R}^{N\times(D+\kappa)},y^{*}_{0}\in\mathbb{R}^{N})$ that is then passed through the NDP de-noising model $\epsilon^{NDP}_{\theta}(x^{*}_{t},y^{*}_{t},t):\mathbb{R}^{N\times(D+\kappa)}\times\mathbb{R}^{N}\to\;\mathbb{R}^{N}$. By constructing the task encoder in this way, we ensure agnosticity to input size. Permutation invariance with respect to the ordering of context points is ensured by aggregating context information through the mean operator $\bar{v}\in\mathbb{R}^{\kappa}$. The target points are kept unchanged by the encoder. This construction is compatible with invariant noise model architectures, as discussed in Section 2.1.
3.3 Objective
@@
3.4 Sampling
Unconditional sampling remains unchanged from the NDP setup described in Section 2, as we use no context points at all. Simply take $y_{T}\sim\mathcal{N}(0,I)$, and then for $t=T,\dots,1$ apply the parametrised kernel. For conditional sampling, we still follow the process for conditional sampling of NDPs described in Section 2, but now using our upgraded de-noising model $\epsilon^{MT}_{\theta}$:
@@ 3.4 Sampling
4 Experiments
Wind farm SCADA data provide a realistic and challenging benchmark for multi-task probabilistic regression. The data are high-frequency, noisy, and high-dimensional, and wind farms consist of multiple turbines that operate under broadly similar environmental conditions while exhibiting turbine-specific behaviour. This setting naturally motivates models that can share information across turbines while adapting to individual dynamics. From a modelling perspective, this application places simultaneous demands on scalability, uncertainty calibration, and cross-task generalisation. Models must handle large datasets spanning multiple years, provide reliable predictive uncertainty, and adapt to turbines with limited historical data. In particular, the ability to condition predictions on a small number of turbine-specific observations at inference time is essential in realistic operational scenarios. These properties make wind farm SCADA data well-suited for evaluating MT-NDPs. In the following experiments, we evaluate model performance in terms of predictive accuracy, uncertainty calibration, and generalisation to unseen turbines under varying numbers of context points.
Dataset. We use SCADA data from the six Senvion MM92 wind turbines at Kelmarsh wind farm in the UK [Plumley2022]. The dataset spans January 2016 to July 2021 at a ten-minute resolution, comprising more than 1.7 million records across 110 variables, including wind speed, wind direction, power output, and component temperatures. Each variable is reported as mean, minimum, maximum, and standard deviation within the ten-minute bins. Further details on SCADA variables and the complete data cleaning procedure are provided in the Appendix C. All preprocessing steps are applied consistently across models.
Baselines. We compare the proposed MT-NDP framework with single-task NDP (ST-NDP) baselines trained independently per wind turbine. All NDP variants share the same diffusion architecture, differing only in whether task-specific information is incorporated. To assess the effect of input dimensionality, ST-NDP models are evaluated using both one-dimensional and five-dimensional feature sets. We additionally compare against GP baselines as a classical probabilistic regression approach. While GPs provide principled uncertainty estimates in small-data regimes, they scale poorly with dataset size and dimensionality, and sparse approximations have been shown to yield underconfident uncertainty estimates on this dataset [Fiocchi_2025]. GP models are trained per turbine using the same data splits and preprocessing as the NDP-based models. Theoretical background and further details for the GP baselines are provided in Appendix A.
@@ 4.1 Training and testing framework
All features are standardised per wind turbine, using the mean and standard deviation from the wind turbine’s training set. This step is necessary for both NDPs and GPs. To ensure fair comparisons, we use the same test functions across all models and context sizes. Context sets of size 0, $25$, and $50$ points are considered. During training, the number of context points in each set is sampled uniformly from the integers 0 to $50$. This setting exposes the model to situations with no context, with few context points, and with many context points. As a result, the model learns both a global prior when no context is provided and task-specific priors when context points are available. This design reflects realistic deployment scenarios, where the amount of available historical data can vary substantially across turbines. Each evaluation uses $30$ test functions of length $100$.
4.2 Model parameters
For all NDP variants, we use a cosine noise schedule [nichol2021improveddenoisingdiffusionprobabilistic] with $T=500$ diffusion steps. The denoising NN has $4$ hidden layers of dimension $64$, with multi-head self-attention using $\kappa=8$ heads. The optimiser is ADAM [Kingma2014ADAM:OPTIMIZATION], following the configuration of [dutordoir2023neural], with a $20$-epoch warmup, $200$-epoch decay, learning rate schedule ($2\times 10^{-5}\to 10^{-3}\to 10^{-5}$), and an EMA weight of $0.995$. For sampling, we generate $200$ reverse diffusion trajectories in the one-dimensional case and $100$ in higher-dimensional settings, balancing predictive sharpness with computational cost.
In the multi-task framework, the task encoder is an MLP with $4$ layers of size $64$, GeLU activations, and an $8$-dimensional output embedding. The mean embedding across all context points forms a task representation vector. Training is performed for $250$ epochs using batches of $32$ function samples of length $100$, with $500$ samples drawn per wind turbine ($5$ tasks/turbines in total), and $100$ randomly shuffled function samples presented per epoch.
Note that the MLP-based encoder does not guarantee invariance to permutations of input features. Enforcing full invariance would require multi-head self-attention blocks, as used in the NDP noise model described in Section 2.1, but would substantially increase model complexity and computational cost. In the wind farm setting, where the ordering of input features can be fixed across training and testing, the lack of invariance is less problematic than it would be in broader meta-learning settings.
@@ 4.3 Model evaluation
To assess point prediction accuracy, we compute the following per-sample losses for each test set: (i) mean absolute error (MAE), defined as $\mathrm{MAE}s=\tfrac{1}{N}\sum_{i=1}^{N}\lvert y_{i}-\hat{y}_{i}\rvert$ and (ii) root mean squared error (RMSE), defined as $\mathrm{RMSE}s=\sqrt{\tfrac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$.
@@
4.4 Illustrative examples
We provide qualitative illustrations of the diffusion dynamics underlying NDPs, complementing the theoretical framework presented in Section 2. The aim is to build intuition for the forward corruption process and for the behaviour of unconditional and conditional sampling during reverse diffusion. We focus on a one-dimensional regression setting for clarity and visualise representative trajectories at different diffusion timesteps. Fig. 1 illustrates the forward diffusion process, where Gaussian noise is gradually added to the outputs using a cosine variance schedule [nichol2021improveddenoisingdiffusionprobabilistic]. As the timestep $t$ increases, the data loses structure and converges towards white noise. This process is applied independently at each input location while the inputs themselves remain fixed.
Unconditional sampling starts from random noise and follows the reverse diffusion process for a fixed number of timesteps. Even without any context information, the resulting samples already resemble the overall data distribution, indicating that the learned prior captures meaningful structure. For wind turbine data, where the relationships between inputs and power output are similar across turbines, the prior alone often produces plausible results.
Conditional sampling incorporates context points at each timestep of the reverse process, guiding the trajectories towards functions that are consistent with the observed data. Initially, when $y^{*}_{t}$ is close to pure noise, the context points already contain useful structure. This pulls the reverse trajectories towards functions consistent with the context set, leading to samples that better fit the true data. Fig. 2 compares unconditional and conditional sampling. Even at early timesteps, the conditional trajectories are visibly influenced by the context points, producing outputs that more closely match the true data throughout the reverse process.
Figure 1: Illustration of the forward diffusion process. Gaussian noise is added incrementally at each timestep under a cosine variance schedule, progressively degrading the structure of the data until it converges to white noise. GP: Gaussian process, CI: Confidence interval.
Figure 2: Comparison of unconditional and conditional reverse diffusion sampling. We sampled 10 reverse trajectories per input location, plotting the mean of these. Conditional sampling provides a closer fit to the true data across all timesteps. GP: Gaussian process.
To illustrate the dynamics more clearly, Fig. 3 depicts the forward and reverse trajectories for a single input point. In the forward process (blue), Gaussian noise gradually corrupts the true value. In the reverse process (red), starting from pure noise, the trajectories converge back to the true value, while uncertainty narrows as the timestep $t$ decreases.
(a) conditional sampling
(b) unconditional sampling
@@ 5 Results
We first compare a one-dimensional GP with a one-dimensional NDP, using only wind speed as a feature. In this setting, wind turbines are nearly indistinguishable, consistent with exploratory analysis (Appendix, Fig. 9(a)). For the GP, we mimic the NDP setup, training on turbines 2266 and testing on turbine $1$. Due to scalability limits, the GP is trained on a subsample of $2000$ points evenly drawn from the five turbines. We compare this GP baseline only to the NDP with no context points, ensuring a fair assessment of prior quality.
This comparison indicates that NDPs can be competitive with GPs on this dataset, motivating their use as a foundation for more expressive models. We then investigate the effect of increasing input dimensionality by comparing one-dimensional and five-dimensional NDPs. The higher-dimensional model additionally incorporates wind direction (encoded via sine and cosine components), nacelle temperature, and transformer temperature. These variables were selected based on exploratory analysis (Fig. 10), which revealed strong and nonlinear relationships with power output. For conciseness, the detailed results for this comparison are reported in the Appendix D.
@@
5.1 GP against NDP
We first compare the one-dimensional NDP without context points against a GP baseline. The effect of context on NDP performance is examined separately in Sections 5.2 and Appendix Section D.
Figure 4: Fit of Gaussian process (GP) (left) and neural diffusion process (NDP) (right) on a test sample of $100$ points from turbine $1$, with 9595% confidence intervals (CI) on normalised wind speed and power output. 1D: One-dimensional.
@@ 5.1 GP against NDP
(a) Mean absolute error (MAE, left) and root mean square error (RMSE, right) distributions for Gaussian process (GP) and neural diffusion process (NDP) over $30$ test samples, each with $100$ points.
(b) Observed vs. theoretical coverage probabilities for Gaussian process (GP) and neural diffusion process (NDP) over $30$ test samples, each with $100$ points.
Figure 5: Comparison of Gaussian process (GP) and neural diffusion process (NDP) on turbine 11 test data. (a) Point prediction accuracy (MAE and RMSE distributions). (b) Predictive uncertainty calibration (coverage probabilities). 1D: One-dimensional, ST: single-task
Considering all $30$ test samples of $100$ points each, both models achieve comparable point prediction accuracy. As shown in Fig. 5(a), the distributions of MAE and RMSE for the GP and NDP are nearly indistinguishable, indicating that both capture the mean power–wind speed relationship effectively. The distinction between the models becomes clear when evaluating predictive uncertainty. Fig. 5(b) shows the empirical coverage against the nominal coverage level. A perfectly calibrated model would follow the diagonal $y=x$ line. While the NDP is not flawless, it consistently lies closer to this ideal than the GP, which suffers from pronounced underconfidence across coverage levels. This highlights a key strength of NDPs: the ability to generate well-calibrated uncertainty estimates without relying on Gaussian assumptions.
@@ 5.2.1 Encoder analysis
These findings have two key implications for evaluation. First, turbine $1$ is not an informative test case for the MT-NDP, as its embeddings remain close to the global prior, and additional context yields little task-specific adaptation. Second, turbine $6$ emerges as the most distinctive and challenging turbine in the wind farm, requiring the model to rely strongly on context information. For this reason, we focus our quantitative evaluation on turbine $6$ in the following sections, as it provides a more discriminative test of the MT-NDP’s ability to adapt beyond the global prior.